Jon has three times as many dimes as he does quarters. He has as many nickels as he has dimes and quarters combined. The total numbers of money he has is 3.00. How many of each coin does he have?: Jon has three times as many dimes as he does quarters. He has as many nickels as he has dimes and quarters combined. The total numbers of money he has is 3.00. How many of each coin does he have?
Answer by orca(336)

Let x be the number of quarters, then the number of dimes is 3x, and the number of nickels is 3x + x = 4x
The value of nickels: 5*4x=20x
The value of dimes: 10*3x=30x
The value of quarters: 25x.
As the total value of money he has is 300 cents, we have
20x + 30x + 25x = 300
Solving for x, we obtain:
75x = 300
x = 4
So he has
4 quarters
3x = 3*4 = 12 dimes
4x = 4*4 = 16 nickels